OpenStudy (here_to_help15):

I will fan and medal Part 1: Suppose you divide a polynomial by a binomial. How do you know if the binomial is a factor of the polynomial? Create a sample problem that has a binomial which IS a factor of the polynomial being divided, and another problem that has a binomial which is NOT a factor of the polynomial being divided.

OpenStudy (here_to_help15):

@campbell_st

OpenStudy (campbell_st):

you use the factor theorem if (x - a) is a factor then f(a) = 0

OpenStudy (here_to_help15):

This is for the first question right?

OpenStudy (campbell_st):

here is a simple example is (x - 3) a factor of x^2 - 8x + 15 well find f(3) f(3) = 9 -24 + 15 f(3) = 0 so (x - 3) is a factor of the polynomial

OpenStudy (campbell_st):

for the not question show (x - 2) isn't a factor f(2) = 4 - 16 + 15 = 3 so not a factor

OpenStudy (here_to_help15):

@campbell_st so is there a way to put this question in a simple answer?

OpenStudy (here_to_help15):

@phi plz help

OpenStudy (campbell_st):

go and google the factor theorem... that is what you are being asked about here is the basics if (x - a) is a factor of a polynomial P(x) then x - a = 0 or x = a is zero of the polynomial. by substituting x = a into the polynomial... and getting P(a) = 0.. you are showing that there is no remainder, when the division occurs so (x - a) is a factor.

OpenStudy (here_to_help15):

But @campbell_st so the question How do you know if the binomial is a factor of the polynomial?is by using the factor theorem?

OpenStudy (here_to_help15):

help @phi?

OpenStudy (phi):

How do you know if the binomial is a factor of the polynomial? Have you learned how to divide a binomial into a polynomial?

OpenStudy (here_to_help15):

i dont know but it would be great if u showed me plz

OpenStudy (phi):

Let me ask it this way: were you supposed to learn how to divide a binomial into a polynomial?

OpenStudy (phi):

or were you taught synthetic division? or were you taught the factor theorem that campbell mentioned? those are 3 ways to tell if a binomial divides evenly into a polynomial

OpenStudy (here_to_help15):

yes i am going to 9th and this is just a review of algebra 1 B

OpenStudy (phi):

which sounds most familiar?

OpenStudy (phi):

Here is how to divide a polynomial by a binomial http://www.khanacademy.org/math/algebra2/polynomial_and_rational/dividing_polynomials/v/dividing-polynomials-1

OpenStudy (phi):

How do you know if the binomial is a factor of the polynomial? one answer is use the factor theorem as posted above. another answer is divide the binomial into the polynomial and see if you get NO remainder

OpenStudy (phi):

Create a sample problem that has a binomial which IS a factor of the polynomial to do this, first write down 2 binomials, and multiply them together (both of your binomials are factors of the polynomial)

OpenStudy (here_to_help15):

So for the first question i could put umm is You can tell if the binomial is a factor of the polynomial is with the Factor Theorem,and Synthetic Division?

OpenStudy (phi):

If you explain the factor theorem or if you say, after using synthetic division you get a remainder of zero

OpenStudy (here_to_help15):

ok so lets just say i was answering this question(which i am) I would start like The way you can tell if the binomial is a factor of the polynomial is if you use the synthetic division and get a remainder of zero.

OpenStudy (here_to_help15):

Then i would give a example

OpenStudy (phi):

that would be good. Can you find a sample problem?

OpenStudy (here_to_help15):

Um i made 1

OpenStudy (here_to_help15):

(2x-7)(5x+3)

OpenStudy (here_to_help15):

so i would use FOIL method right?

OpenStudy (here_to_help15):

@phi i gave a sample would this be good?

OpenStudy (here_to_help15):